[Brahms-l] Shift report 20040225 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Wed Feb 25 2004 - 00:01:10 EST
------------------------------------------------------------
Date: Wed Feb 25 00:01:10 2004, Shift: 16:00-24:00
Supervisor: BSN, with: PJ
Log book#15 pages 66 - 66

Shift Summary:
------------------------------------------------------------
1. Problems storing first half of the shift.

2.
        FS 8deg. 1/2A    MRS 40deg. 1/2A    MRS 40deg. full A 

10584   66k              130k               0
10585   70k              212k               0
10586   11k               57k               0
10587   Still running       

3. Bad trip alarm for T2 in run 10584, but it didn't seem to shut down the HV (don't know why). But a new run was started not too long after anyway.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Feb 25 00:01:28 2004

This archive was generated by hypermail 2.1.8 : Wed Feb 25 2004 - 00:01:44 EST